Government funding: The following table presents the budget (in millions of dollars) for selected organizations that received U.S. government funding for arts and culture in both 2004 and 2006. Use a TI-84 calculator to answer the following.www.own.connectmath.com s - Playlist Push MATHI6-0621-FA19 McOwl Education Campus McGraw-Hill Connect Question 2 of 5 (1 poin
Answer #1

let x represent 2004 values and y represents 2006 values.

Sum of X = 1026
Sum of Y = 1139
Mean X = 205.2
Mean Y = 227.8
X-M₂ Y-M 283.8 171.8 -70.2 -188.2 -197.2 289.2 232.2 -85.8 -209. -225.8 (X - Mx)2 80542.44 29515.24 4928.04 35419.24 38887.84Sum of squares (55x) = 189292.8 Sum of products (SP) = 212002.2 Regression Equation = ĝ= bX+ a b= SPISSx= 212002.2/189292.8 =